The perimeter of a rectangle is 50 ft. Describe the possible length of a side if the area of the rectangle is not to exceed 114 square feet.

L * W > 114
2(L * W) = 50
L + W = 25
-W -W
L = 25 – W

Answer by Alan3354(69443)

The perimeter of a rectangle is 50 ft. Describe the possible length of a side if the area of the rectangle is not to exceed 114 square feet.

L * W > 114
2(L * W) = 50
L + W = 25
-W -W
L = 25 – W
-----------------
It's L*W <= 114 for "not to exceed"
Sub for L
W*(25-W) <= 114
W^2 - 25W + 114 <=0
(W - 6)*(W - 19) <= 0
--> 6 or 19
W = 6, L = 19 or
W = 19, L = 6
--------------
W <= 6 for area <= 114
If one side is <= 6, the area is <= 114.